OpenStudy (haleyelizabeth2017):

What is the equation, in standard form, of a parabola that contains the following points? (–2, –20), (0, –4), (4, –20)

OpenStudy (haleyelizabeth2017):

I know \[y=ax^2+bx+c\] is the standard form for the equation.....

OpenStudy (haleyelizabeth2017):

I don't know why it has the diamonds but the points are: (-2, -20), (0, -4) and (4, -20)

OpenStudy (here_to_help15):

is this word question?

OpenStudy (noelgreco):

You need to solve three simultaneous equations. The first is\[-20 = (-2)^{2}a+(-2)b+c\] All I did was substitute y with -20, and x with -2 from the first point (-2,-20). Do the same with the next two points and solve the three equations for the three coefficients a, b, and c.

OpenStudy (haleyelizabeth2017):

its a workpad :'( so I haveta show all work.....:( I forgot how to do this :/

OpenStudy (haleyelizabeth2017):

okay so: \[-20=(-2)^2a+(-2)b+c ; (-4)=(0)^2a+(0)b+c ; -20=(4)^2a+(4)b+c\] right?

OpenStudy (here_to_help15):

ok hold on

OpenStudy (haleyelizabeth2017):

Sorry it took so long....I had to go do something for my mom....:/

OpenStudy (here_to_help15):

y = ax^2 + bx + c -20 = 4a - 2b + c -4 = c -20 = 16a + 4b + c 4a - 2b + c = 16a + 4b + c 4a - 2b = 16a + 4b -6b = 12a b = -2a -20 = 4a - 2b + c -20 = 4a + 4a - 4 -16 = 8a -2 = a b = 4 c = -4 y = -2x^2 + 4x - 4 y = -2 * (x^2 - 2x) - 4 y = -2 * (x^2 - 2x + 1 - 1) - 4 y = -2 * (x^2 - 2x + 1) + 2 - 4 y = -2 * (x + 1)^2 - 2

OpenStudy (here_to_help15):

lol i answered this on yahoo

OpenStudy (here_to_help15):

with a help of friend

OpenStudy (here_to_help15):

@haleyelizabeth2017

OpenStudy (haleyelizabeth2017):

I have to have the steps :( Like I need to explain them....I think I can figure it out though

OpenStudy (here_to_help15):

i did give you steps lol

OpenStudy (haleyelizabeth2017):

I meant I needa make the explanations haha :P

OpenStudy (here_to_help15):

-_-

OpenStudy (here_to_help15):

hey

OpenStudy (haleyelizabeth2017):

yo

OpenStudy (here_to_help15):

you aint no nig ga so stop acting like 1 :P

OpenStudy (haleyelizabeth2017):

wtf???

OpenStudy (here_to_help15):

:O IM CALLING A MOD HAHAHA JK :P

OpenStudy (haleyelizabeth2017):

you the one sayin bad things!

OpenStudy (here_to_help15):

O NO YOUR THE ONE MADE ME CRY SNIFF SNIFF YOu HUT MY FEELINGS WITH UR BAD WORDS

OpenStudy (here_to_help15):

hurt

OpenStudy (haleyelizabeth2017):

how????????????? I said yo den wtf.....I never said it. It could mean when to freak :P

OpenStudy (here_to_help15):

MHMM LIAR :P

OpenStudy (here_to_help15):

MOD MOD MOD

OpenStudy (here_to_help15):

lol you are making me die i just cant stop smiling

OpenStudy (haleyelizabeth2017):

shut up. buhbye!!! im not talking to you for rest of day

OpenStudy (here_to_help15):

NO NO NO SORRY NO PLEASE :'( but fine!IIF YOU WANT LEAVE ME AGAIN ITS NOTHING NEW!!!

OpenStudy (haleyelizabeth2017):

:P

OpenStudy (here_to_help15):

i am trying so hard to be mad i just cant :/

OpenStudy (here_to_help15):

but seriously sorry :(

OpenStudy (haleyelizabeth2017):

haha

OpenStudy (here_to_help15):

WHATS SO FRUCKING FUNNY

OpenStudy (haleyelizabeth2017):

your face :P jk jk

OpenStudy (here_to_help15):

>:( FU CK OS

OpenStudy (here_to_help15):

IMA GO GET BANNED

OpenStudy (here_to_help15):

Ill include you in chat :P

OpenStudy (haleyelizabeth2017):

NOOOOOOOOOO Ducky you :P jk jk

OpenStudy (here_to_help15):

lol

OpenStudy (here_to_help15):

ok ducky me :(

OpenStudy (here_to_help15):

k bye :'(

OpenStudy (here_to_help15):

I SAID BYE!

OpenStudy (haleyelizabeth2017):

huh?

OpenStudy (here_to_help15):

geez la fwucking weeze

OpenStudy (haleyelizabeth2017):

loser buhbye!!! :P

OpenStudy (here_to_help15):

thats it!!

OpenStudy (here_to_help15):

Do ya even like me or do i bother you?tell me huh?

OpenStudy (here_to_help15):

ya thats what i thought

OpenStudy (here_to_help15):

:/ bye haley and to save you from being pissed ill leave you alone!

OpenStudy (haleyelizabeth2017):

huh? i was reading! calm down!

OpenStudy (here_to_help15):

Ok just answer the question if im bothering you just tell me

OpenStudy (haleyelizabeth2017):

nope

OpenStudy (here_to_help15):

you sure :/

OpenStudy (haleyelizabeth2017):

yup

OpenStudy (here_to_help15):

ok but just block me if i annoy u ok?

OpenStudy (haleyelizabeth2017):

okay :P

OpenStudy (here_to_help15):

but lemme know before u do lol

OpenStudy (here_to_help15):

so i can change ya mind

OpenStudy (haleyelizabeth2017):

heehee

OpenStudy (here_to_help15):

lol

OpenStudy (haleyelizabeth2017):

That isn't right.....:( I don't think